answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What are 3 probabilities of getting heads when you flip a quarter?

There is the probability of 1/2 if it is a fair coin.

There is the probability of 1 if it is a double-headed coin.

There is the probability of 0 if it is a double-tailed coin.

When a straight thumbed homozygous person has offspring with a hitchhiker's thumbed homozygous person the offspring will have which trait?

Straight thumbs, hitch hiker's thumb is recessive...so the F1 generation in this example would be heterozygous and have the dominant (straight thumbed) phenotype.

What is the probability of spinning a blue and then a red or a yellow there are four colors on the board?

If there are four colors on a spinner, then the probability of spinning one particular color is 1 in 4, or 0.25. Also, the probability of spinning one of two particular colors is 2 in 4, or 0.5.

Combining these two "unrelated" events simply requires multiplication. The probability, then, of spinning one particular color on one spin, and then spinning one of two particular colors on the next spin is (1 in 4) times (2 in 4), or 2 in 16, or 0.125.

What is the probability of drawing a red card or a spade in two draws without replacing the card?

There are 26 red cards and 13 spades in a standard deck of 52 cards. The probability of drawing a red card or a spade in one draw is, therefore, 39 in 52. If you draw twocards, and the first is not red or spade, then the probability on the second draw is 39 in 51, otherwise it is 38 in 51.

Combining these two probabilities is easy. Just turn the problem around, and ask what is the probability of drawing two clubs? The answer is (13 in 52) times (12 in 51), which is 156 in 2652, or 1 in 17. Flip that answer over by subtracting it from 1, and you get a probability of drawing a red card or a spade in two draws of 16 in 17, or about 0.9412.

What is the probability of drawing a blue card replacing it and then drawing a blue card?

The answer depends on how many cards there are, what colours, how many of each colour, whether or not the cards are in random order. And since you have not bothered to share any of this crucial information, I cannot provide a more useful answer.

The answer depends on how many cards there are, what colours, how many of each colour, whether or not the cards are in random order. And since you have not bothered to share any of this crucial information, I cannot provide a more useful answer.

The answer depends on how many cards there are, what colours, how many of each colour, whether or not the cards are in random order. And since you have not bothered to share any of this crucial information, I cannot provide a more useful answer.

The answer depends on how many cards there are, what colours, how many of each colour, whether or not the cards are in random order. And since you have not bothered to share any of this crucial information, I cannot provide a more useful answer.

What does the expression mutually exclusive mean?

Mutually exclusive means if one thing is true the other must be false and vice versa.
If A is true, B is false. If B is true, A is false.

For instance,
The ball was totally red.
The ball was totally blue.
These are mutually exclusive because the ball can only be one or the other.

The ball was red.
The ball was blue.
These are NOT mutually exclusive because the ball could also be red AND blue.

How powerful is the Jack of spades?

Power level of a Jack of Spades is OVER 9000

What are your chances of winning at roulette?

Your chances of winning in roulette are 1 in 38. If you play on black or red, you will double your money about 45 % of the time. Some people will do better and will come out ahead. For everyone who does, someone will do worse. If you play long enough, the house will always win.

What does Numismatics mean?

yes numismatics is the science of all numbers basically numismatics is study of coins from ancient history till the present it helps us to reconstruct history

What does inclusivity mean?

Diversity and inclusion exists when members of an organization act in a manner that recognizes and respects individual similarities and differences such that employees feel they and their work are valued and meaningfully important to the success of the company. Increasing workforce diversification challenges all members of a company attempting to satisfy the individual value. Differences in age, race, gender, religion, and sexual orientation to name but only a few, influence what makes individuals feel valued. Studies have shown the degree to which an individual feels valued by his/her superiors and peers in large part defines his/her work engagement and the value offered to the company. The challenge is a circular one best solved by fostering an organized culture that respects and achieves to embrace diversity and inclusion. But what makes an employee feel valued? Studies show that what makes us feel valued differs depending on our particular position and background. Often, these studies help the needs, values, and attitudes of individuals sharing a common background to provide broad insight to what makes members of these groups seek and make them feel valued. While not perfect, these studies provide a starting point for becoming aware of and understanding individual preferences; awareness and understanding that is becoming increasingly important given the rapidly changing workforce demographic. We can make our community more inclusive by uniting behind the decision: * Demonstrate the group's commitment to equality. * Bring to light different skills, experiences and perspectives, and new solutions. * Increase the company to new knowledge, networks and connections, helping to open doors and increase the community. * Bring special characteristics to the company- for example, young adults often help to start enthusiasm and can also bring with them new perspectives. * Improve communication with other co-workers. * Help to keep the company fresh. We need to examine the company. Changing the culture of a company through more inclusion will often require a change in the way things are done. Ask questions like: * What has your company done to encourage inclusion in the past? * What worked? What wasn't so successful? * Where does the company stand now with regards to inclusion? * Where would your company like to be with regards to inclusion in 12 months time? * How will you achieve this? * Know why you are doing it. A person may feel they have to work twice as hard and be twice as successful to achieve the same respect and recognition others might. Try to avoid stereotyping. Observe, listen to and take cues from the people you are trying to include and take this into consideration in your interaction with them. "Treat others as you would like to be treated."

What are the steps to doing probability?

The steps are:

1. TRY IT

2. Ask questions

3. figure out the problem

4. SHOW ALL YOUR WORK

5. turn it in

Hope this was help. Probally was not.

What is the entire group that the researcher is studying?

The entire group that the researcher is interested in is called the population or the target population.

Sample selected in such a way that each member of the population has an equal mark?

(Simple) Random sampling users a sample size of n in which every sample of size n has an equal change of being chosen.

How many cards are in a pinochle deck?

48 cards. In double Pinochle (the only kind I have ever played), you use two Pinochle decks, and take out all the 9s. The relative rank of the cards is switched from most card games somewhat as well... A is still high, but then 10 comes next, followed by K, Q, and J (and then 9 if you use it).

How many 6 in a deck of cards?

4 cards but if you want the probability of getting a 6 in pack of cards it is 4/52=1/13

A fair die is rolled 180 times what is the probability of rolling a six?

The probability of rolling a six is 1 out of 6, or 1/6. Now, perhaps your question is:

If I roll a die 180 times, what is the probability of rolling a six at least once. This is the same as rolling a die 180 times and never once rolling a six. The probability is (5/6)180 which is 5.59 x 10-15.

Two cards are dealt from a standard 52 card deck What is the probability that both cards are from the suit of diamonds?

There is one chance in four of the first card being a diamond (since there are four suits). There is one chance in four of the second card being a diamond. The chances of both being a diamond is 1/4*1/4 = 1/16, or thus one chance in 16